Can you expand: 1/(a+b) to something like c/d + e/f ?
Thank you!

Answer by wgunther(43)

c=0, d=anything but 0, e=1, f=a+b.
c=2, d=a+b, e=-1, f=a+b
c=1, d=1, e=1-a-b, f=a+b
Not very helpful I'm sure, but algebraically 1/(a+b) cannot be expanded into anything more helpful then it is. There are an infinite number of rational sums however.